1.     A random sample of n = 100 measurements is obtained from a population with 𝜇 = 55 and 𝜎 = 20.  Describe the sampling distribution for the sample means by computing the 𝜇𝑥̅ and 𝜎𝑥̅


1
Expert's answer
2022-03-04T05:58:15-0500

Since the sample size n=100 is statistically considered to be large (n>30), we assume that the mean of the sampling distribution of  is equal to the population mean, μ.

Therefore;

 𝜇𝑥̅  = μ = 55

Also, we estimate the sample standard deviation (𝜎𝑥̅.) using the formula:


𝜎𝑥̅. = "\frac{\sigma}{\sqrt{n}}"

where "\sigma" = population standard deviation


Given n=100 and "\sigma" = 20

Then;

𝜎𝑥̅ = "\frac{\sigma}{\sqrt{n}}" = "\frac{20}{\sqrt{100}}" = 2


Answer: The sampling distribution of the sample means has 𝜇𝑥̅ = 55 and 𝜎𝑥̅ =2